# Intuitionistic NF set theory

### v5

@@ 8. Addition
$Nc(SSC(b\cup\{c\})=Nc(SSC(b))+Nc(SSC(b)).$
Proof. Define
$R:=\{x\cup\{c\}:x\in SSC(b)\}.$
The definition can be rewritten in stratified form, so $R$ can be defined in INF. Define $f:x\mapsto x\cup\{c\}$, which can also be defined in INF:
$f:=\{\langle x,x\cup\{c\}\rangle:x\in SSC(b)\}.$
The formula is stratified, since all the occurrences of $x$ can be given index 0, and $\{c\}$ and $SSC(b)$ are just parameters. Then $f:SSC(b)\to R$ is a similarity. (We omit the 150 steps required to prove that.)
We first note that if $x\in SSC(b\cup\{c\})$ and $c\in x$, then $x=(x-c)\cup\{c\}$, since $x$ is finite and therefore has decidable equality. Similarly $b\cup\{c\}$ has decidable equality, so every $x\in SSC(b\cup\{c\})$ either contains $c$ or not. If $c\in x$ then $x\in R$. If $c\not\in x$ then $x\in SSC(b)$. Therefore
$SSC(b\cup\{c\})=SSC(b)\cup R$
$SSC(b)\sim R$   since $f:SSC(b)\to R$ is a similarity
$Nc(SSC(b))=Nc(R)$   by Lemma 4.12
$SSC(b)\cap R=\Lambda$   since $c\not\in b$
$SSC(b)\in{\sf FINITE}$   by Lemma 3.17
@@
$Nc(SSC(b\cup\{c\}))=Nc(SSC(b))+Nc(R)$ by Lemma 8.15
$Nc(SSC(b\cup\{c\}))=Nc(SSC(b))+Nc(SSC(b))$ since $Nc(SSC(b))=Nc(R)$
That completes the proof of the lemma.
@@
Proof. Suppose $p\in{\mathbb{F}}$ and $2^{p^{+}}\in{\mathbb{F}}$. Then
@@
$USC(a)\in p^{+}$   for some $a\in p$, by definition of exponentiation
@@
$u\sim USC(a)$   by Lemma 4.9, since both are in $p^{+}$
$w\in USC(a)$   for some $w$, by Lemma 6.20
$c\in a\land w=\{c\}$   for some $c$, by definition of $USC(a)$
$USC(a)\in{\sf FINITE}$   by Lemma 4.4
@@
$a=b\cup\{c\}$   since $a\in{\sf DECIDABLE}$
$USC(a)=USC(b)\cup\{\{c\}\}$   by Lemma 6.15
$SSC(b\cup\{c\})\in 2^{p^{+}}$   by definition of exponentiation
$Nc(SSC(b\cup\{c\})=2^{p^{+}}$   by Lemma 4.11
(63) $\ \ Nc(SSC(b\cup\{c\})=Nc(SSC(b))+Nc(SSC(b))$ by Lemma 8.17
$USC(a)\in{\sf DECIDABLE}$   by Lemma 3.3
$USC(b)=USC(a)-\{\{c\}\}$   since $USC(a)\in{\sf DECIDABLE}$
$USC(b)\in p$   by Lemma 5.10
$SSC(b)\in 2^{p}$   by the definition of exponentiation
$SSC(b)\in Nc(SSC(b))$   by Lemma 4.11
$SSC(b)\in{\sf FINITE}$   by Lemma 3.17
$Nc(SSC(b))\in{\mathbb{F}}$   by Lemma 4.21
$Nc(SSC(b))=2^{p}$   by Lemma 5.7
Then $2^{p^{+}}=2^{p}+2^{p}$ as desired, by (63). That completes the proof of the lemma.
@@
Proof. Left to right. We have
@@
Right to left. Suppose $m={\sf zero}$. Then $2^{m}=2^{\sf zero}={\sf one}$, by Lemma 7.6. That completes the proof of the lemma.
@@
Proof. We prove by induction on $n$ that for $n\in{\mathbb{F}}$ with $2^{n}$ inhabited, we have
(64) $\exists u\,(u\in 2^{n})\ {\rightarrow}\ \forall m\in{\mathbb{F}}\,(2^{n}=2^{m}\ {\rightarrow}\ n=m)$
@@
Case 3, $n=r$. Then $2^{n}=2^{r}$. Substituting $2^{n}$ for $2^{r}$ in the identity $2^{r}+2^{r}=2^{r}+2^{r}$, we have $2^{n}+2^{n}=2^{r}+2^{r}$. Then $2^{n^{+}}=2^{r^{+}}=2^{m}$ as desired. That completes the induction step. That completes the proof of the lemma.
@@
Proof. Suppose $m<n$ and $2^{n}$ is inhabited. Then
$m\leq n$   by the definition of $<$
@@
$2^{m}<2^{n}$   by the definition of $<$
That completes the proof of the lemma.
@@
Proof. By induction on $q$. The formula is stratified, giving all variables index 0.
Base case, $p\leq{\sf zero}{\ \leftrightarrow\ }\exists k\in{\mathbb{F}},p+k={\sf zero}$. Left to right: Suppose $p\leq{\sf zero}$. Then $p={\sf zero}\ \lor\ p<{\sf zero}$, by Lemma 5.20. But $p\not<{\sf zero}$ by Lemma 5.28. Hence $p={\sf zero}$. Then $p+k={\sf zero}+k={\sf zero}$ by Lemma 8.3. . Right to left. Suppose $p+k={\sf zero}.$ Then by the definition of addition, there exists sets $a\in p$ and $b\in k$ such that $a\cup b\in{\sf zero}$. By definition of ${\sf zero}$, ${\sf zero}=\{\Lambda\}$, so $a\cup b=\Lambda$. Then $a=\Lambda$. Then $\Lambda\in p$ and $\Lambda\in{\sf zero}$. Then by Lemma 5.7, $p={\sf zero}$. That completes the base case.
@@
Right to left. Suppose $k\in{\mathbb{F}}$ and $p+k=q^{+}$. We have to show $p\leq q^{+}$. By definition of addition, there exist $a$ and $b$ with $a\in p$ and $b\in k$, and $a\cap b=\Lambda$ and $a\cup b\in q^{+}$. Then $a$ is a separable subset of $a\cup b$, so $p\leq q^{+}$ by the definition of $\leq$. That completes the induction step. That completes the proof of the lemma.
@@
Proof. Suppose $p,q\in{\mathbb{F}}$ and $p+q\in{\mathbb{F}}$. We have
@@
$\Lambda\in{\sf zero}$   by the definition of ${\sf zero}$
$\Lambda\subseteq u\ \land\ u=\Lambda\cup(u-\Lambda)$   by the definitions of subset and difference
@@
That is the first assertion of the lemma. By Lemma 8.3, we have $p+q=q+p$, so $q+p\in{\mathbb{F}}$ and as above we have $q\leq q+p$. Therefore also $q\leq p+q$. That completes the proof of the lemma.
@@
Proof. By induction on $p$. For the base case, there is nothing to prove. For the induction step, assume $p^{+}$ is inhabited and $2^{p^{+}}\in{\mathbb{F}}$ and $p^{+}\neq{\sf zero}$ and $p^{+}\neq{\sf one}$. We have to prove
(65) $p^{++}<2^{p^{+}}$
@@
(66) $p\neq{\sf zero}$   since $p^{+}\neq{\sf one}$ by hypothesis
@@
(68) $2^{p}\in{\mathbb{F}}$   by Lemma 7.5, since it is inhabited
@@
But that is (65), the desired goal. That completes the induction step. That completes the proof of the lemma.
9. Definition of multiplication
Specker did not make any use of multiplication. If one could manage to prove that ${\mathbb{F}}$ is infinite, one would need multiplication to interpret HA in INF. But without knowing that ${\mathbb{F}}$ is infinite, there are technical difficulties with multiplication. Some care is required to make sure that the equations for multiplication work without assuming ${\mathbb{F}}$ is finite; the equations must have the property that if one side is in ${\mathbb{F}}$, so is the other side. That is, if one side “overflows”, so does the other side. To arrange this, we must first ensure that addition has the same property. This ultimately goes back to the theorem that successor never takes the value zero, not just on an integer argument but on any argument whatever. We carried out those details (and they can still be found in earlier versions of this paper on ArXiv), but we have not included them here.
10. Results about ${\mathbb{T}}$
@@
${\mathbb{T}}(\kappa)$ $=$ $\{u:\exists x\,(x\in\kappa\ \land\ u\sim USC(x))\}$
@@
$x\in\kappa{\ \leftrightarrow\ }USC(x)\in{\mathbb{T}}\kappa.$
Proof. Left to right:
@@
$USC(x)\sim USC(x)$   by Lemma 2.11
$USC(x)\in{\mathbb{T}}(\kappa)$   by Definition 10.1
@@
$USC(x)\in{\mathbb{T}}\kappa$   by hypothesis
$\exists z\,(z\in\kappa\ \land\ USC(z)\sim USC(x))$   by definition of ${\mathbb{T}}$
@@
That completes the right-to-left direction. That completes the proof.
@@
If $\kappa\in{\mathbb{F}}$ then for every $x\in\kappa$, ${\mathbb{T}}(\kappa)=Nc(USC(x))$.
Proof. Suppose $\kappa\in{\mathbb{F}}$. Then $\kappa$ is inhabited, by Corollary 4.7. Let $x\in\kappa$. Then
@@
$USC(x)\in{\sf FINITE}$   by Lemma 3.10
$Nc(USC(x))\in{\mathbb{F}}$   by Lemma 4.21
$USC(x)\in{\mathbb{T}}(\kappa)$   by Lemma 10.2
$USC(x)\sim USC(x)$   by Lemma 2.11
$USC(x)\in Nc(USC(x))$   by Definition 4.10
@@
(69) $\forall u\,(u\in{\mathbb{T}}(\kappa){\ \leftrightarrow\ }u\in Nc(USC(x)))$
Left to right: Suppose $u\in{\mathbb{T}}(\kappa)$. By definition of ${\mathbb{T}}$, there exists $w\in\kappa$ with $u\sim USC(w)$. Then
@@
$USC(w)\sim USC(x)$   by Lemma 6.7
$u\sim USC(x)$   by Lemma 2.11 (transitivity of $\sim$), since $u\sim USC(w)$
@@
Right to left: Suppose $u\in Nc(USC(x))$. Then $u\sim USC(x)$. Since $x\in\kappa$, we have $u\in{\mathbb{T}}(\kappa)$ by the definition of ${\mathbb{T}}$. That completes the proof of the lemma.
@@
If $\kappa\in{\mathbb{F}}$ and $x\in\kappa$ then $\kappa=Nc(x)$.
Proof. Let $\kappa\in{\mathbb{F}}$ and $x\in\kappa$. By extensionality, it suffices to prove that for all $u$,
$u\in\kappa{\ \leftrightarrow\ }u\in Nc(x).$
@@
$u\in Nc(x)$   by Definition 4.10
Right to left: Suppose $u\in Nc(x)$. Then
@@
If $Nc(x)\in{\mathbb{F}}$, then ${\mathbb{T}}(Nc(x))=Nc(USC(x))$.
Proof. By Lemma 10.3, with $\kappa=Nc(x)$.
@@
Proof. Let $m\in{\mathbb{F}}$. By Corollary 4.7, $m$ is inhabited. Let $a\in m$. Then
$USC(a)\in{\mathbb{T}}m$   by Lemma 10.2
@@
$USC(a)\in{\sf FINITE}$   by Lemma 3.10
$Nc(USC(a))\in{\mathbb{F}}$   by Lemma 4.21
@@
Every singleton has cardinal ${\sf one}$. That is, $\forall x\,(Nc(\{x\})={\sf one}).$
Proof. By definition, ${\sf one}={\sf zero}^{+}$ and ${\sf zero}=\{\Lambda\}$. Then the members of ${\sf one}$ are sets of the form $\Lambda\cup\{r\}$, by the definition of successor. But $\Lambda\cup\{r\}=\{r\}$. Hence the members of ${\sf one}$ are exactly the unit classes. Let $x$ be given; then by definition of $Nc$, $Nc(\{x\})$ contains exactly the sets similar to $\{x\}$. By Lemma 6.3, that is exactly the unit classes. Hence Nc({x}) and ${\sf one}$ have the same members, namely all unit classes. By extensionality, $Nc(\{x\})={\sf one}$. That completes the proof.
@@
Proof. Since $m^{+}$ is inhabited, there is an $x\in m$ and $a\not\in x$ (so $x\cup\{a\}\in m^{+}$). Then
@@
Proof. We have $USC(\Lambda)=\Lambda$ as there are no singleton subsets of $\Lambda$. Since ${\sf zero}=Nc(\Lambda)$, by Lemma 10.5 we have ${\mathbb{T}}({\sf zero})=Nc(USC(\Lambda))=Nc(\Lambda)={\sf zero}$. That completes the proof that ${\mathbb{T}}({\sf zero})={\sf zero}.$
@@
$\{\Lambda\}\in{\sf one}$   by definition of ${\sf one}$
${\mathbb{T}}({\sf one})=Nc(USC(\{\Lambda\})$   by Lemma 10.3
${\mathbb{T}}({\sf one})=Nc(\{\{\Lambda\}\})$   since $USC(\{\Lambda\})=\{\{\Lambda\}\}$
$Nc(\{\{\Lambda\}\})={\sf one}$   by Lemma 10.7
@@
Proof. The formula in the lemma is stratified, with the relation $<$ occurring as a parameter. Therefore we can prove by induction that for $n\in{\mathbb{F}}$,
@@
and only one of the three disjuncts holds. Therefore it suffices to rule out the first two disjuncts, as the third is the desired conclusion. By Lemma 5.33, the first one is impossible. We turn to the second. Suppose ${\mathbb{T}}m={\sf zero}$. Since $m\in{\mathbb{F}}$, by Lemma 4.7 we have $a\in m$ for some $a$. Then $USC(a)\in{\mathbb{T}}m$, by definition of ${\mathbb{T}}$. Since ${\mathbb{T}}m={\sf zero}$, we have $USC(a)\in{\sf zero}$. Since ${\sf zero}=\{\Lambda\}$, we have $USC(a)=\Lambda$. Then $a=\Lambda$. Since $\Lambda\in{\sf zero}$, by Lemma 5.7 and the fact that $a\in m$, we have $m={\sf zero}$. But that contradicts the assumption ${\sf zero}<m$, by Lemma 5.33. That completes the base case.
@@
That completes the induction step. That completes the proof of the lemma.
@@
Proof. By induction on $m$ we prove
(70) $\forall n\in{\mathbb{F}}\,(n+m\in{\mathbb{F}}\ {\rightarrow}\ {\mathbb{T}}(n+m)={\mathbb{T}}n+{\mathbb{T}}m).$
@@
That is the desired goal. That completes the induction step. That completes the proof of the lemma.
@@
Proof. Let $m\in{\mathbb{F}}$. Then
@@
$USC(u)\in{\mathbb{T}}m$   by Definition 10.1
$SSC(u)\in 2^{{\mathbb{T}}m}$   by the definition of exponentiation
That completes the proof of the lemma.

### v6

@@ 8 Addition
$|\mathcal{P}_{s}(b\cup\{c\})|=|\mathcal{P}_{s}(b)|+|\mathcal{P}_{s}(b)|.$
Proof.
Define
$R:=\{x\cup\{c\}:x\in\mathcal{P}_{s}(b)\}.$
The definition can be rewritten in stratified form, so $R$ can be defined in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. Define $f:x\mapsto x\cup\{c\}$, which can also be defined in $\mathit{i}\mkern 0.4mu\mathrm{NF}$:
$f:=\{\langle x,x\cup\{c\}\rangle:x\in\mathcal{P}_{s}(b)\}.$
The formula is stratified, since all the occurrences of $x$ can be given index 0, and $\{c\}$ and $\mathcal{P}_{s}(b)$ are just parameters. Then $f:\mathcal{P}_{s}(b)\to R$ is a similarity. (We omit the 150 steps required to prove that.)
We first note that if $x\in\mathcal{P}_{s}(b\cup\{c\})$ and $c\in x$, then $x=(x-c)\cup\{c\}$, since $x$ is finite and therefore has decidable equality. Similarly $b\cup\{c\}$ has decidable equality, so every $x\in\mathcal{P}_{s}(b\cup\{c\})$ either contains $c$ or not. If $c\in x$ then $x\in R$. If $c\not\in x$ then $x\in\mathcal{P}_{s}(b)$. Therefore
$\mathcal{P}_{s}(b\cup\{c\})=\mathcal{P}_{s}(b)\cup R$
$\mathcal{P}_{s}(b)\sim R$   since $f:\mathcal{P}_{s}(b)\to R$ is a similarity
$|\mathcal{P}_{s}(b|)=|R|$   by Lemma 4.12
$\mathcal{P}_{s}(b)\cap R=\emptyset$   since $c\not\in b$
$\mathcal{P}_{s}(b)\in{\sf FINITE}$   by Lemma 3.17
@@
$|\mathcal{P}_{s}(b\cup\{c\}|)=|\mathcal{P}_{s}(b|)+|R|$ by Lemma 8.15
$|\mathcal{P}_{s}(b\cup\{c\}|)=|\mathcal{P}_{s}(b|)+|\mathcal{P}_{s}(b|)$ since $|\mathcal{P}_{s}(b|)=|R|$
@@
Proof.
Suppose $p\in{\mathbb{F}}$ and $2^{p^{+}}\in{\mathbb{F}}$. Then
@@
$\mathcal{P}_{1}(a)\in p^{+}$   for some $a\in p$, by definition of exponentiation
@@
$u\sim\mathcal{P}_{1}(a)$   by Lemma 4.9, since both are in $p^{+}$
$w\in\mathcal{P}_{1}(a)$   for some $w$, by Lemma 6.20
$c\in a\land w=\{c\}$   for some $c$, by definition of $\mathcal{P}_{1}(a)$
$\mathcal{P}_{1}(a)\in{\sf FINITE}$   by Lemma 4.4
@@
$a=b\cup\{c\}$   since $a\in{\sf DECIDABLE}$ (64)
$\mathcal{P}_{1}(a)=\mathcal{P}_{1}(b)\cup\{\{c\}\}$   by Lemma 6.15
$\mathcal{P}_{s}(b\cup\{c\})\in 2^{p^{+}}$   by definition of exponentiation
$|\mathcal{P}_{s}(b\cup\{c\}|=2^{p^{+}}$   by Lemma 4.11
$|\mathcal{P}_{s}(b\cup\{c\})|=|\mathcal{P}_{s}(b)|+|\mathcal{P}_{s}(b)|$ by Lemma 8.17 (65)
$\mathcal{P}_{1}(a)\in{\sf DECIDABLE}$   by Lemma 3.3
$\mathcal{P}_{1}(b)=\mathcal{P}_{1}(a)-\{\{c\}\}$   since $\mathcal{P}_{1}(a)\in{\sf DECIDABLE}$
$\mathcal{P}_{1}(b)\in p$   by Lemma 5.9
$\mathcal{P}_{s}(b)\in 2^{p}$   by the definition of exponentiation
$\mathcal{P}_{s}(b)\in|\mathcal{P}_{s}(b|)$   by Lemma 4.11
$\mathcal{P}_{s}(b)\in{\sf FINITE}$   by Lemma 3.17
$|\mathcal{P}_{s}(b)|\in{\mathbb{F}}$   by Lemma 4.21
$|\mathcal{P}_{s}(b)|=2^{p}$   by Lemma 4.24
Then $2^{p^{+}}=2^{p}+2^{p}$ as desired, by (65). ∎
@@
Proof.
Left to right. We have
@@
Right to left. Suppose $m={\sf zero}$. Then $2^{m}=2^{\sf zero}={\sf one}$, by Lemma 7.6. ∎
@@
Proof.
We prove by induction on $n$ that for $n\in{\mathbb{F}}$ with $2^{n}$ inhabited, we have
$\exists u\,(u\in 2^{n})\ {\rightarrow}\ \forall m\in{\mathbb{F}}\,(2^{n}=2^{m}\ {\rightarrow}\ n=m)$ (66)
@@
Case 3, $n=r$. Then $2^{n}=2^{r}$. Substituting $2^{n}$ for $2^{r}$ in the identity $2^{r}+2^{r}=2^{r}+2^{r}$, we have $2^{n}+2^{n}=2^{r}+2^{r}$. Then $2^{n^{+}}=2^{r^{+}}=2^{m}$ as desired. That completes the induction step. ∎
@@
Proof.
Suppose $m<n$ and $2^{n}$ is inhabited. Then
$m\leq n$   by the definition of <<
@@
$2^{m}<2^{n}$   by the definition of <<
@@
Proof.
By induction on $q$. The formula is stratified, giving all variables index 0.
Base case, $p\leq{\sf zero}{\ \leftrightarrow\ }\exists k\in{\mathbb{F}},p+k={\sf zero}$. Left to right: Suppose $p\leq{\sf zero}$. Then $p={\sf zero}\ \lor\ p<{\sf zero}$, by Lemma 5.19. But $p\not<{\sf zero}$ by Lemma 5.29. Hence $p={\sf zero}$. Then $p+k={\sf zero}+k={\sf zero}$ by Lemma 8.3.
. Right to left. Suppose $p+k={\sf zero}$. Then by the definition of addition, there exists sets $a\in p$ and $b\in k$ such that $a\cup b\in{\sf zero}$. By definition of $\sf zero$}, $\sf zero$$\{\emptyset\}$, so $a\cup b=\emptyset$. Then $a=\emptyset$. Then $\emptyset\in p$ and $\emptyset\in{\sf zero}$. Then by Lemma 4.24, $p={\sf zero}$. That completes the base case.
@@
Right to left. Suppose $k\in{\mathbb{F}}$ and $p+k=q^{+}$. We have to show $p\leq q^{+}$. By definition of addition, there exist $a$ and $b$ with $a\in p$ and $b\in k$, and $a\cap b=\emptyset$ and $a\cup b\in q^{+}$. Then $a$ is a separable subset of $a\cup b$, so $p\leq q^{+}$ by the definition of $\leq$. That completes the induction step. ∎
@@
Proof.
Suppose $p,q\in{\mathbb{F}}$ and $p+q\in{\mathbb{F}}$. We have
@@
$\emptyset\in{\sf zero}$   by the definition of $\sf zero$}
$\emptyset\subseteq u\ \land\ u=\emptyset\cup(u-\emptyset)$   by the definitions of subset and difference
@@
That is the first assertion of the lemma. By Lemma 8.3, we have $p+q=q+p$, so $q+p\in{\mathbb{F}}$ and as above we have $q\leq q+p$. Therefore also $q\leq p+q$. ∎
@@
Proof.
By induction on $p$. For the base case, there is nothing to prove. For the induction step, assume $p^{+}$ is inhabited and $2^{p^{+}}\in{\mathbb{F}}$ and $p^{+}\neq{\sf zero}$ and $p^{+}\neq{\sf one}$. We have to prove
$p^{++}<2^{p^{+}}$ (67)
@@
$p\neq{\sf zero}$   since $p^{+}\neq{\sf one}$ by hypothesis (68)
@@
$2^{p}\in{\mathbb{F}}$   by Lemma 7.5, since it is inhabited (70)
@@
But that is (67), the desired goal. That completes the induction step. ∎
Lemma 8.25.
Let $q\in{\mathbb{F}}$. Then for all $n\in{\mathbb{F}}$ and $p\in{\mathbb{F}}$,
[…]
Remark. This lemma links addition and order. It probably can be proved directly from the definitions of addition and order, but here we prove it by induction. Nevertheless we do have to use the definition of addition directly at one of the steps.
Proof.
By induction on $q$, which is legal since the formula is stratified. The formula to be proved includes the quantifiers on $n$ and $p$.
Base case. There is nothing to prove because of the hypothesis $q\neq{\sf zero}$.
Induction step. Suppose $n=p+q^{+}$ and ${\sf zero}<q^{+}$. As usual in induction proofs, we also assume $q^{+}$ is inhabited. Then
$n=p^{+}+q$   by Lemma 8.2
$q<{\sf zero}\ \lor\ q={\sf zero}\ \lor\ {\sf zero}<q$   by Theorem 5.16
Case 1, $q<{\sf zero}$ is impossible, by Lemma 5.29.
Case 2, $q={\sf zero}$. Then $q^{+}={\sf one}$ so $n=p+{\sf one}=p^{+}$. Then $p<n$ by Lemma 5.37.
Case 3, ${\sf zero}<q$. Then
$n\mbox{\ is inhabited}$   by Lemma 4.7
$p^{+}\mbox{\ is inhabited}$   by the definition of addition, since $n\ p^{+}q$
$p^{+}<n$   by the induction hypothesis, since $n=p^{+}+q$
$p<p^{+}$   by Lemma 5.37
$p<n$   by transitivity
That completes the induction step. ∎
9 Definition of multiplication
Specker did not make any use of multiplication. If one could manage to prove that $\mathbb{F}$} is infinite, one would need multiplication to interpret HA in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. But without knowing that $\mathbb{F}$} is infinite, there are technical difficulties with multiplication. Some care is required to make sure that the equations for multiplication work without assuming $\mathbb{F}$} is finite; the equations must have the property that if one side is in $\mathbb{F}$}, so is the other side. That is, if one side “overflows”, so does the other side. To arrange this, we must first ensure that addition has the same property. This ultimately goes back to the theorem that successor never takes the value zero, not just on an integer argument but on any argument whatever. We carried out those details (and they can still be found in earlier versions of this paper on ArXiv), but we have not included them here.
10 Results about $\mathbb{T}$}
@@
${\mathbb{T}}(\kappa)$ $=$ $\{u:\exists x\,(x\in\kappa\ \land\ u\sim\mathcal{P}_{1}(x))\}$
@@
$x\in\kappa{\ \leftrightarrow\ }\mathcal{P}_{1}(x)\in{\mathbb{T}}\kappa.$
Proof.
Left to right:
@@
$\mathcal{P}_{1}(x)\sim\mathcal{P}_{1}(x)$   by Lemma 2.11
$\mathcal{P}_{1}(x)\in{\mathbb{T}}(\kappa)$   by Definition 10.1
@@
$\mathcal{P}_{1}(x)\in{\mathbb{T}}\kappa$   by hypothesis
$\exists z\,(z\in\kappa\ \land\ \mathcal{P}_{1}(z)\sim\mathcal{P}_{1}(x))$   by definition of $\mathbb{T}$}
@@
That completes the right-to-left direction. ∎
@@
If $\kappa\in{\mathbb{F}}$ then for every $x\in\kappa$, ${\mathbb{T}}(\kappa)=|\mathcal{P}_{1}(x|)$.
Proof.
Suppose $\kappa\in{\mathbb{F}}$. Then $\kappa$ is inhabited, by Corollary 4.7. Let $x\in\kappa$. Then
@@
$\mathcal{P}_{1}(x)\in{\sf FINITE}$   by Lemma 3.10
$|\mathcal{P}_{1}(x|)\in{\mathbb{F}}$   by Lemma 4.21
$\mathcal{P}_{1}(x)\in{\mathbb{T}}(\kappa)$   by Lemma 10.2
$\mathcal{P}_{1}(x)\sim\mathcal{P}_{1}(x)$   by Lemma 2.11
$\mathcal{P}_{1}(x)\in|\mathcal{P}_{1}(x|)$   by Definition 4.10
@@
$\forall u\,(u\in{\mathbb{T}}(\kappa){\ \leftrightarrow\ }u\in|\mathcal{P}_{1}(x|))$ (71)
Left to right: Suppose $u\in{\mathbb{T}}(\kappa)$. By definition of $\mathbb{T}$}, there exists $w\in\kappa$ with $u\sim\mathcal{P}_{1}(w)$. Then
@@
$\mathcal{P}_{1}(w)\sim\mathcal{P}_{1}(x)$   by Lemma 6.7
$u\sim\mathcal{P}_{1}(x)$   by Lemma 2.11 (transitivity of $\sim$), since $u\sim\mathcal{P}_{1}(w)$
@@
Right to left: Suppose $u\in|\mathcal{P}_{1}(x|)$. Then $u\sim\mathcal{P}_{1}(x)$. Since $x\in\kappa$, we have $u\in{\mathbb{T}}(\kappa)$ by the definition of $\mathbb{T}$}. ∎
@@
If $\kappa\in{\mathbb{F}}$ and $x\in\kappa$ then $\kappa=|x|$.
Proof.
Let $\kappa\in{\mathbb{F}}$ and $x\in\kappa$. By extensionality, it suffices to prove that for all $u$,
$u\in\kappa{\ \leftrightarrow\ }u\in|x|.$
@@
$u\in|x|$   by Definition 4.10
Right to left: Suppose $u\in|x|$. Then
@@
If $|x|\in{\mathbb{F}}$, then ${\mathbb{T}}(|x|)=|\mathcal{P}_{1}(x|)$.
Proof.
By Lemma 10.3, with $\kappa=|x|$. ∎
@@
Proof.
Let $m\in{\mathbb{F}}$. By Corollary 4.7, $m$ is inhabited. Let $a\in m$. Then
$\mathcal{P}_{1}(a)\in{\mathbb{T}}m$   by Lemma 10.2
@@
$\mathcal{P}_{1}(a)\in{\sf FINITE}$   by Lemma 3.10
$|\mathcal{P}_{1}(a|)\in{\mathbb{F}}$   by Lemma 4.21
@@
Every singleton has cardinal $\sf one$}. That is, $\forall x\,(|\{x\}|={\sf one})$.
Proof.
By definition, ${\sf one}={\sf zero}^{+}$ and $\sf zero$$\{\emptyset\}$. Then the members of $\sf one$} are sets of the form $\emptyset\cup\{r\}$, by the definition of successor. But $\emptyset\cup\{r\}=\{r\}$. Hence the members of $\sf one$} are exactly the unit classes. Let $x$ be given; then by definition of $|\{x\}|$, $|\{x\}|$ contains exactly the sets similar to $\{x\}$. By Lemma 6.3, that is exactly the unit classes. Hence $|\{x\}|$ and $\sf one$} have the same members, namely all unit classes. By extensionality, $|\{x\}|={\sf one}$. ∎
@@
Proof.
Since $m^{+}$ is inhabited, there is an $x\in m$ and $a\not\in x$ (so $x\cup\{a\}\in m^{+}$). Then
@@
Proof.
We have $\mathcal{P}_{1}(\emptyset)=\emptyset$ as there are no singleton subsets of $\emptyset$. Since ${\sf zero}=|\emptyset|$, by Lemma 10.5 we have ${\mathbb{T}}({\sf zero})=|\mathcal{P}_{1}(\emptyset|)=|\emptyset|={\sf zero}$. ∎
@@
$\{\emptyset\}\in{\sf one}$   by definition of $\sf one$}
${\mathbb{T}}({\sf one})=|\mathcal{P}_{1}(\{\emptyset\}|$   by Lemma 10.3
${\mathbb{T}}({\sf one})=|\{\{\emptyset\}\}|$   since $\mathcal{P}_{1}(\{\emptyset\})=\{\{\emptyset\}\}$
$|\{\{\emptyset\}\}|={\sf one}$   by Lemma 10.7
@@
Proof.
The formula in the lemma is stratified, with the relation << occurring as a parameter. Therefore we can prove by induction that for $n\in{\mathbb{F}}$,
@@
and only one of the three disjuncts holds. Therefore it suffices to rule out the first two disjuncts, as the third is the desired conclusion. By Lemma 5.34, the first one is impossible. We turn to the second. Suppose ${\mathbb{T}}m={\sf zero}$. Since $m\in{\mathbb{F}}$, by Lemma 4.7 we have $a\in m$ for some $a$. Then $\mathcal{P}_{1}(a)\in{\mathbb{T}}m$, by definition of $\mathbb{T}$}. Since ${\mathbb{T}}m={\sf zero}$, we have $\mathcal{P}_{1}(a)\in{\sf zero}$. Since $\sf zero$$\{\emptyset\}$, we have $\mathcal{P}_{1}(a)=\emptyset$. Then $a=\emptyset$. Since $\emptyset\in{\sf zero}$, by Lemma 4.24 and the fact that $a\in m$, we have $m={\sf zero}$. But that contradicts the assumption ${\sf zero}<m$, by Lemma 5.34. That completes the base case.
@@
That completes the induction step. ∎
@@
Proof.
By induction on $m$ we prove
$\forall n\in{\mathbb{F}}\,(n+m\in{\mathbb{F}}\ {\rightarrow}\ {\mathbb{T}}(n+m)={\mathbb{T}}n+{\mathbb{T}}m).$ (72)
@@
That is the desired goal. That completes the induction step. ∎
@@
Proof.
Let $m\in{\mathbb{F}}$. Then
@@
$\mathcal{P}_{1}(u)\in{\mathbb{T}}m$   by Definition 10.1
$\mathcal{P}_{s}(u)\in 2^{{\mathbb{T}}m}$   by the definition of exponentiation